OpenStudy (anonymous):

help? -1/2 sqrt 5 + 2 sqrt 125

OpenStudy (anonymous):

\[-\frac{ 1 }{ 2 } \sqrt{5} + 2 \sqrt{125}\]

OpenStudy (cwrw238):

hint: √125 = 5 √5

OpenStudy (anonymous):

doesnt 5 x 25 = 125 though?

OpenStudy (cwrw238):

yes and √( 5 X 25) = √5 X √25 = 5√5

OpenStudy (anonymous):

oh okay i understand that

OpenStudy (cwrw238):

so it becomes -0.5 √5 + 2*5√5 can you finish this off?

OpenStudy (anonymous):

i think so hold on

OpenStudy (cwrw238):

now treat the √5 as if it it were a variable like x

OpenStudy (anonymous):

yeah i dont think i got it right. i got 47.5

OpenStudy (cwrw238):

its - √5 + 10 √5 = 9.5 √5

OpenStudy (cwrw238):

thats ok for your answer theres no need to work this out any further do u see what i mean about treating the surd (i.e. sqrt5) as an 'x'.

OpenStudy (anonymous):

yeah i believe so

OpenStudy (cwrw238):

* - 0.5 √5

OpenStudy (cwrw238):

-0.5x + 10x = 9.5x

OpenStudy (cwrw238):

ok?